evaluate 25 1/2. answer is 2 to the power 4 how do i get to to it?
Answer by rwm(914)

2^4=16
2^4=2^2*2^2=4*4=16
25^1/2=5 since 25^(1/2)=sqrt(25)
5^2=25
I suspect something is wrong here. You copied the wrong answer or the book is wrong.